A current of 75.7 mA exists in a metal wire. How

Need help with these
Answer #1

003) Use equation,

I= dq/dt =(ne)/dt

75.7*10^-3=( n*1.602*10^-19)/(9.4*60)         => n=2.67*10^20

004)

Resistance for 0.17m , R = (2.39*10^-5)*0.17 = 4.063*10^-6 ohm

V=I*R = 122A* 4.063*10^-6 ohm = 0.000496 V = 496 microV